OpenStudy (anonymous):

Simplify the expression by writing as a single term. 81x^4y/9x7^z^3

OpenStudy (anonymous):

@johnweldon1993 can u help me plz

OpenStudy (johnweldon1993):

\[\large \frac{81x^4y}{9x^7z^3}\] This can be broken up like this... \[\large \frac{81}{9} \times \frac{x^4}{x^7} \times\frac{y}{1}\times\frac{1}{z^3}\] can you simplify any of these?